# Phase separation of chemokinetic active particles

### v1

@@ A.2 Case of the AMR
Figure S2: Phase diagram of the quorum-sensing ABPs in the BMR at $\lambda=0.5$ and $D_{\mathrm{c}}=10$. The system size is set to be $L=400$. The LSA prediction for the H–O boundary is indicated by the dotted curve. The snapshots of particle configurations are taken at $t=2\times 10^{3}$, with the values of $\zeta$ and $n_{0}$ given by the coordinates of each snapshot’s center.
In Fig. S2, we show the patterns to which the clusters formed by the quorum-sensing ABPs in Fig. 8(a) evolve if we run the simulations to $t=2\times 10^{3}$. We observe that the small particle clusters merge via the Ostwald process to form large-scale clusters, in agreement with the hydrodynamic prediction. Also note that the particle clusters tend to become elongated as $n_{0}$ becomes smaller, which is also observed in the hydrodynamic model.
@@
Figure S3: Time-variability of the quorum-sensing ABPs in the AMR for $\tau=200$ and $D_{\mathrm{c}}=0.25$ as $\lambda$ and $L_{\mathrm{win}}$ are varied. The other parameters are identical to those used in Fig. 8(b).
Appendix B Extra results for quorum-sensing ABPs
@@
In Fig. S3, we show how the time-variability $\mathcal{V}_{\tau}$ depends on the window size $L_{\mathrm{win}}$ for various values of $\lambda$ along the $D_{\mathrm{c}}=0.25$ line of Fig. 8(b). We note that the value of $L_{\mathrm{win}}$ at which $\mathcal{V}_{\tau}$ is maximized can be regarded as the length scale of oscillating patterns. Then, Fig. S3 indicates that oscillations occur at a shorter length scale with smaller amplitudes as $\lambda$ grows. This is consistent with the system crossing over from the O phase at small $\lambda$ to the H phase at large $\lambda$.

### v2

@@ Appendix E Extra results for quorum-sensing active particles
Figure A5: Phase diagram of the quorum-sensing active Brownian particles in the Basal Metabolic Regime. We use $\lambda=0.5$ and $D_{\mathrm{c}}=10$, with the system size fixed at $L=400$. The LSA prediction for the H–O boundary is indicated by the dotted curve. The snapshots of particle configurations are taken at $t=2\times 10^{3}$, with the values of $\zeta$ and $n_{0}$ given by the coordinates of each snapshot’s center.
@@
Figure A6: Time-variability of the quorum-sensing active Brownian particles in the Active Metabolic Regime. We fix $\tau=200$ and $D_{\mathrm{c}}=0.25$ as $\lambda$ and $L_{\mathrm{win}}$ are varied. The other parameters are identical to those used in Fig. 6(b) in the main text.
In this note, we present extra results for quorum-sensing ABPs. In Fig. 6, we show the patterns to which the clusters formed by the quorum-sensing ABPs in Fig. Hydrodynamic model(a) in the main text in evolve if we run the simulations to $t=2\times 10^{3}$. We observe that the small particle clusters merge via the Ostwald process to form large-scale clusters, in agreement with the hydrodynamic prediction. Also note that the particle clusters tend to become elongated as $n_{0}$ becomes smaller, which is also observed in the hydrodynamic model.
@@
In Fig. A6, we show how the time-variability $\mathcal{V}_{\tau}$ depends on the window size $L_{\mathrm{win}}$ for various values of $\lambda$ along the $D_{\mathrm{c}}=0.25$ line of Fig. 6(b) in the main text. We note that the value of $L_{\mathrm{win}}$ at which $\mathcal{V}_{\tau}$ is maximized can be regarded as the length scale of oscillating patterns. Then, Fig. A6 indicates that oscillations occur at a shorter length scale with smaller amplitudes as $\lambda$ grows. This is consistent with the system crossing over from the O phase at small $\lambda$ to the H phase at large $\lambda$.